OpenStudy (anonymous):

Solve by the substitution 3x-5y=-10 x=-3-2y type an ordered pair

OpenStudy (anonymous):

Since you know what x is, put that in for x in the first equation.

OpenStudy (diyadiya):

3(-3-2y)-5y=-10 -9-6y-5y=-10 -11y=-1 y=1/11

OpenStudy (diyadiya):

by substituting -3-2y in place of x

OpenStudy (anonymous):

3(-3-2y) - 5y = -10 (distribute) -9 - 6y - 5y = -10 -11y - 9 = -10 -11y = -1 y = 1/11 (are you sure your problem is correct, teachers usually give you "nicer" answers. but I will continue)

OpenStudy (anonymous):

x = -3 - 2y x = -3 - 2(1/11) x = -3 - 2/11 x = -3 2/11

OpenStudy (anonymous):

so the orderd pair is a fraction?

OpenStudy (anonymous):

(-3 2/11, 1/11)

OpenStudy (anonymous):

wow, that's a pain to graph!

OpenStudy (anonymous):

I wonder if the two equations are either entered correctly or there was a mistake in the problem... Teachers usually don't do this.

OpenStudy (anonymous):

let me check , I'll be back in a minute.Thank you for your patience.

OpenStudy (anonymous):

I checked, the real question was,"What is the solution to the system, type an ordered pair." so all I need is the solution in coordinate terms

OpenStudy (anonymous):

OK.. so (-3 2/11, 1/11)

OpenStudy (anonymous):

it appears your right

OpenStudy (anonymous):

thanks